The Bouchard Company's EPS was $6.96 in 2016, up from $3.34 in 2011. The company pays out 45% of its earnings as dividends, and its common stock sells for $33.

  1. Calculate the past growth rate in earnings. (Hint: This is a 5-year growth period.) Round your answer to two decimal places.
    %
  2. The last dividend was D0 = 0.45($6.96) = $3.13. Calculate the nextexpected dividend, D1, assuming that the past growth rate continues. Do not round off intermediate calculations. Round your answer to the nearest cent.
    $
  3. What is Bouchard's cost of retained earnings, rs? Round your answer to two decimal places. Do not round your intermediate calculations.
    %

please Help!Thank you!
Answer #1


solution: L FV=PV*(1+r)^n 6.96=3.34(1+r)^5 (1+r)^3=6.96/3.34 (1+r)^3=2.08383 1+r=2.08383^(1/5) (1+r)=1.158171 r=1.158171-1 r=